OpenStudy (anonymous):

The difference of 3 times a number and -2 is 4 :) Please help and explain Im not good at word problems.

OpenStudy (anonymous):

let the number be x. 3 time implies 3x. difference implies 3x - (-2) = 4. Now solve

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

can you help m solve?

OpenStudy (anonymous):

this is quite easy. 3x + 2 = 4 3x = 4-2 3x =2 x = 2/3

OpenStudy (anonymous):

is that x= 2div 3?

OpenStudy (anonymous):

yh

OpenStudy (anonymous):

do I solve that or that is the answer ? :)

OpenStudy (anonymous):

that's the final resutls. \[x = \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

ohh thank you so much :) tho I have more of those:P